OpenStudy (anonymous):

what is the solution of the system 5x - y = -7, 3x - y = -2

OpenStudy (anonymous):

@texaschic101 @charlotte123

OpenStudy (anonymous):

would that be infinity solution

OpenStudy (charlotte123):

Indeed :)

OpenStudy (anonymous):

@charlotte123 can you help me with a couple more questions

OpenStudy (charlotte123):

Sure!

OpenStudy (anonymous):

what is the solution of the system 5x - y = -7, 3x - y = -2

OpenStudy (anonymous):

@texaschic101 help

OpenStudy (radar):

-y = -7 -5x -y = -2-3x -y=-y then -7-5x=-2-3x -2x = -5 x=5/2 or 2 1/2

OpenStudy (anonymous):

is it positive or negative

OpenStudy (radar):

It is the results of a division of two negative numbers or like signs. What do you think?

OpenStudy (anonymous):

positive